OpenStudy (anonymous):

Solve (8x + 1) + (3x – 2) = 32.

OpenStudy (gorv):

11x-1=32 11x=33 x=3

OpenStudy (anonymous):

8x + 1 +3x - 2 = 32 combine like terms 11x - 1 = 32 11x - 1 + 1 = 32+1 11x = 33 sovle for x